OpenStudy (anonymous):

IN NEED OF SOMEONE TO LOOK OVER MY WORK........ A machinist creates a washer by drilling a hole through the center of a circular piece of metal. If the piece of metal has a radius of x + 10 and the hole has a radius of x + 6, what is the area of the washer?

OpenStudy (anonymous):

help boost your brain by playing xbox :DD

OpenStudy (anonymous):

I would love to but I have to get my homework done first.. ): @sourwing

OpenStudy (anonymous):

@mangorox Could you please help me on this problem?

OpenStudy (anonymous):

how would you start this off? just tell me how you would solve it

OpenStudy (anonymous):

A=Pie(x=10^2)- Pie(x+6)^2 A=Pie(x^2+20x+100-x^2-12x-36) A=Pie(8x+64) A=8Pie(x+8) is this correct?

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

\[\Large A = 8\pi(x+8)\] is correct You can also write it as \[\Large A = 8\pi x + 64\pi\]

OpenStudy (anonymous):

But they don't look anything like the answers unless im not looking at them right..

OpenStudy (anonymous):

Ohhhh! I see it now! Thank you!